VIBRATIONAL CORRECTIONS TO MOLECULAR PROPERTIES: SECOND-ORDER VIBRATIONAL PERTURBATION THEORY VS VARIATIONAL COMPUTATIONS

\underline{MICHAEL E. HARDING}, JUANA VÁZQUEZ, JOHN F. STANTON, Institute for Theoretical Chemistry, Department of Chemistry and Biochemistry, University of Texas at Austin, Austin, TX 78712 , USA; GREGOR DIEZEMANN, and JÜRGEN GAUSS, Institut für Physikalische Chemie, Universität Mainz, Jakob-Welder-Weg 11, D-55128 Mainz, Germany.

For a small set of linear and non-linear molecules, a detailed comparison of two different procedures for predicting vibrationally averaged molecular properties, i.e., second-order vibrational perturbation theory (VPT2) and a variational approach, is carried out. Results for vibrational corrections to dipole and quadrupole moments, nuclear quadrupole moments, static electric-dipole polarizabilities, NMR chemical shielding tensors, nuclear spin-rotation tensors, magnetizabilities, and rotational g-tensors are reported.